How many diopters are there for for a converging lens with a focal length of 0.4m?

−2.5
−0.4
+0.4
+2.5
1
Expert's answer
2018-03-31T10:05:08-0400
A lens with a focal length of f=+0.4 m (converging lens) has a power
P=1/f=1/(+0.4m)=+2.5 diopters.